Can you help me solve these problems step by step
3^2 X 3^x = 3^8
15^12/5^12 = x^12
5^4 X 6^4 = x^4

Found 2 solutions by Fombitz, ewatrrr:
Answer by Fombitz(13828)
Answer by ewatrrr(10682)


Hi,
If I am understanding the Question properly:
solving
3^2 * 3^x = 3^8
3^x = 3^8/3^2
3^x = 3^6
x = 6
15^12/5^12 = x^12
(3*5)^12/5^12 = x^12
3^12*5^12/5^12 = x^12
3^12 = x^12
3 = x
5^4 * 6^4 = x^4
(5*6)^4 = x^4
30 = x